# Inheritune: Training Smaller Yet More Attentive Language Models

### v2

@@ B.1 Main Results in Low Data Regime
B.2 Additional analysis with larger reference LMs and 50B data
@@
Table 9: MMLU (5-shot) scores of Our-1.5B small base LM derived using 1B data for multiple data repetition–10 epochs and 20 epochs compared to the same model trained without data repetition for 10B and 20B fresh tokens. We derive all the variants of Our-1.5B small base using $\mathsf{Inheritune}$ with OpenLLaMA-3B as reference model. The models featured in this table correspond to those discussed in Figures 12 and 13.
@@
B.3 Implications of Low Data Regime
@@
Appendix C Implementation Details
C.1 Training details of GPT-2 models
@@
C.2 Training details of 1.5B OpenLLaMA model

### v3

@@ E.1 Main Results in Low Data Regime
E.2 Additional analysis with larger reference LMs and 50B data
@@
Table 10: MMLU (5-shot) scores of Our-1.5B small base LM derived using 1B data for multiple data repetition–10 epochs and 20 epochs compared to the same model trained without data repetition for 10B and 20B fresh tokens. We derive all the variants of Our-1.5B small base using $\mathsf{Inheritune}$ with OpenLLaMA-3B as reference model. The models featured in this table correspond to those discussed in Figures 21 and 22.
@@
E.3 Implications of Low Data Regime
@@
Appendix F Implementation Details
F.1 Training details of GPT-2 models
@@
F.2 Training details of 1.5B OpenLLaMA model